OpenStudy (anonymous):

Which expression is equivalent to –3x + 9?

OpenStudy (anonymous):

What are you expressions that you have to compare it to?

OpenStudy (anonymous):

A. –3(x + 9) B. –3(x – 3) C. –3(x + 12) D. 3(–x – 3)

OpenStudy (anonymous):

Okay, first, solve the given equation: -3x + 9 = 0 -3x = -9 x = 3

OpenStudy (anonymous):

im really bad a solving equations :/

OpenStudy (anonymous):

Now plug in 3 to the given answers and see if it = 0. I'll do A for you. -3(x + 9) = 0 -3((3) + 9) = 0 -3(12) = 0 -36 = 0 False. So a is not correct.

OpenStudy (anonymous):

Do you understand what I'm doing?

OpenStudy (anonymous):

to be honest no....i get really confused

OpenStudy (anonymous):

That's okay. Think of x as a question mark: x = ? 2 + ? = 3 Do you know what that is?

OpenStudy (anonymous):

not really....i just dont understand

OpenStudy (anonymous):

Okay, solve this: 1 + ____ = 2

OpenStudy (anonymous):

1 plus what equals 2

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Yeah, so: 1 + x = 2 x is just a place holder, or an unknown number. x + 6 = 7 x + 4 = 5 For both of those equations, x = 1. Do you understand how x works now?

OpenStudy (dan815):

sky hello

OpenStudy (anonymous):

Sky, what grade are you in?

OpenStudy (dan815):

We are factoring out the common multiple -3x+9 remember hat 3*3=9 so -3x+3*3<---- as you can see 3 or -3 is common in both terms so we can "factor" it -3(x-3)<---- you can do a check to see if this is right by multiplying through the terms by -3

OpenStudy (anonymous):

dan that's too hard for her.

OpenStudy (anonymous):

i feel stupid.

OpenStudy (anonymous):

im a freshman

OpenStudy (dan815):

You're not stupid